OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (4, -1) and perpendicular to the line whose equation is 2x - y - 7 = 0 y = -1/2x + 1 y = 1/2x - 1 y = 2x + 1

OpenStudy (anonymous):

C

sam (.sam.):

Nope

OpenStudy (anonymous):

could you explain ho to work it out??

OpenStudy (anonymous):

how*

sam (.sam.):

What is the gradient perpendicular to 2x - y - 7 = 0 ?

OpenStudy (anonymous):

i have no idea ma"am i havent learned that termonology yet

OpenStudy (anonymous):

1x-2y-7=0??

sam (.sam.):

You can find the gradient from 2x - y - 7 = 0, 2x - y - 7 = 0 y = 2x - 7 From y = mx + b m is the gradient, 2 --------------------------- Then to find the perpendicular gradient, use \[m_1m_2=-1 \\ 2m_2=-1 \\ m_2=-\frac{1}{2}\] --------------------------- Now you have a gradient and coordinate (4,-1) You can now use the point slope form to finish it up \[y-y_1=m(x-x_1) \\ y+1=-\frac{1}{2}(x-4)\]\[...\]